# Mu-MASS Collaboration

### v1

@@ I Introduction
As a purely leptonic atom, Muonium ($\mathrm{M}$}), a hydrogen-like bound state of a positive muon ($\mathrm{\mu^{+}}$) and an electron ($e^{-}$), is free from hadronic structure effects and finite-size contributions. The absence of nuclear structure makes it a uniquely clean system for high-precision tests of bound-state quantum electrodynamics (QED)[1, 2, 3, Patkos:2024lqf, 4]. Its relatively long lifetime of $2.2,\mu\mathrm{s}$, in contrast to the much shorter one of positronium, allows for spectroscopic measurements with reduced natural line broadening. Precision spectroscopy of Muonium provides the most accurate determinations of the muon mass and magnetic moment [5, 6], thereby serving as an essential input to Standard Model precision tests. Beyond QED studies, Muonium also provides a sensitive probe of new physics, including searches for Lorentz and CPT violation[7, 8], couplings to dark sectors [9, 10], and possible new or exotic long-range forces [11, 12, 13].
@@ IV Results and Discussion
Figure 6: Comparison of ${\mathrm{H}}\ 2$ and ${\mathrm{M}}\ 5$ resonance scans. The dashed gray lines indicate the underlying lineshapes for visualization only. Left: Muonium scan ${\mathrm{M}}\ 5$ in red circles with the corresponding best fit from the simulation represented by blue triangles: $f_{\mathrm{{\mathrm{M}}5}}=-6.0\pm 9.1\,\mathrm{MHz}$, $\chi_{\mathrm{min,{\mathrm{M}}5}}^{2}=7.7$. Right: Hydrogen scan ${\mathrm{H}}\ 2$ in red circles with the corresponding best fit from the simulation represented by blue triangles: $f_{\mathrm{{\mathrm{H}}2}}=2.1\pm 9.1\,\mathrm{MHz}$, $\chi_{\mathrm{min,{\mathrm{H}}2}}^{2}=2.6$.

### v3

@@ I Introduction
As a purely leptonic atom, Muonium ($\mathrm{M}$}), a hydrogen-like bound state of a positive muon ($\mathrm{\mu^{+}}$) and an electron ($e^{-}$), is free from hadronic structure effects and finite-size contributions. The absence of nuclear structure makes it a uniquely clean system for high-precision tests of bound-state quantum electrodynamics (QED)[1, 2, 3, 4, 5]. Its relatively long lifetime of $2.2,\mu\mathrm{s}$, in contrast to the much shorter one of positronium, allows for spectroscopic measurements with reduced natural line broadening. Precision spectroscopy of Muonium provides the most accurate determinations of the muon mass and magnetic moment [6, 7], thereby serving as an essential input to Standard Model precision tests. Beyond QED studies, Muonium also provides a sensitive probe of new physics, including searches for Lorentz and CPT violation[8, 9], couplings to dark sectors [10, 11], and possible new or exotic long-range forces [12, 13, 14, 15].
@@ IV Results and Discussion
Figure 6: Comparison of ${\mathrm{H}}\ 2$ and ${\mathrm{M}}\ 5$ resonance scans. The dashed gray lines indicate the underlying lineshapes for visualization only. Left: Hydrogen scan ${\mathrm{H}}\ 2$ in red circles with the corresponding best fit from the simulation represented by blue triangles: $f_{\mathrm{{\mathrm{H}}2}}=2.1\pm 9.1\,\mathrm{MHz}$, $\chi_{\mathrm{min,{\mathrm{H}}2}}^{2}=2.6$. Right: Muonium scan ${\mathrm{M}}\ 5$ in red circles with the corresponding best fit from the simulation represented by blue triangles: $f_{\mathrm{{\mathrm{M}}5}}=-6.0\pm 9.1\,\mathrm{MHz}$, $\chi_{\mathrm{min,{\mathrm{M}}5}}^{2}=7.7$.